OpenStudy (anonymous):

If the dimensions of a rectangle become one-fifth of the original dimensions, then area of the new rectangle will be _______of the area of the original rectangle.

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

@asnaseer

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

For a rectangle with dimensions a, b, area is given by x: x = a * b When we reduce those dimensions to 1/5th in our new rectangle for which area is y: y = (a/5) * (b/5) y = (a*b)/(5*5) y = (a*b)/25 Substituting: y = x/25 Hope this helps.

OpenStudy (anonymous):

The answer is 25? How would that go in to the question :P?

OpenStudy (anonymous):

"one twenty-fifth"

OpenStudy (anonymous):

ty :D but the thing is that they don't explain is are you suppose to say it one twenty-fifth or 1/25 ;(

OpenStudy (anonymous):

it probably doesn't matter, you'll probably get the marks either way :)